# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 3.1.4 Synchrotron Amplitude
Deriving a full-sky template for synchrotron emission in total intensity is challenging, both due to the paucity of low-frequency surveys with full sky coverage and the difficulty in disentangling synchrotron emission from free-free emission and AME at frequencies above $\sim$10 GHz. We follow Thorne et al. (2017) in basing our template on the Haslam 408 MHz survey (Haslam et al., 1982) and in particular the reprocessing of the original maps by Remazeilles et al. (2015). We rescale the 408 MHz map to 23 GHz assuming a sky-constant scaling of the brightness temperature with $\nu^{-3.1}$. Finally, we smooth the resulting map with a $2^{\circ}$ Gaussian beam.
@@
A key limitation of the current synchrotron total intensity template is that, despite the point source masking, we still find excess power from Galactic point sources at low Galactic latitudes near the Galactic center. This excess power can then leak into our final polarization maps through the polarization fraction tensor transformations (Equations (3.1.1) and (3.1.1)). Our models are unaffected by this issue at $\ell\lesssim 3000$ and outside the Galactic plane. Addressing small-scale structures in bright regions will be deferred to a future release.
@@
3.2 Small-Scale Fluctuations in Spectral Parameters
3.2.1 Methods Overview
@@
As there is no sense of polarization in the spectral parameter maps, we do not utilize the polarization fraction tensor framework, but rather work with standard power spectra. Given a template map $\mathcal{T}$ of some spectral parameter $X$, we first fit the power spectrum of $\mathcal{T}$ with a power law $\mathcal{D}_{\ell}^{XX}\propto\ell^{\alpha}$ over a range of scales $\ell_{0}<\ell<\ell_{1}$ where it is well-measured. We then generate a map of small-scale fluctuations using the power law fit. We multiply the resulting map by a modulation map, in analogy with the amplitude modulations described in Section 3.1.2. Finally, we combine the template map with the synthesized small scales using the filter function in Equation 10. As with the amplitude parameters, we use the input template maps for the spectral parameters as-is in the inner Galactic plane, defined by the Planck GAL097 mask.
@@
3.2.2 Dust Spectral Parameters
As described in Section 2.2.1, the dust emission in all models developed here is governed by the spectral parameters $\beta_{d}$ and $T_{d}$. For the $\beta_{d}$ and $T_{d}$ template maps, we employ the $\beta_{d}$ and $T_{d}$ maps [^10] COM_CompMap_Dust-GNILC-Model-Spectral-Index_2048_R2.00.fits, COM_CompMap_Dust-GNILC-Model-Temperature_2048_R2.00.fits derived from 2015 Planck GNILC component separation analysis (Planck Collaboration Int. XLVIII, 2016). These maps have lower CIB contamination than the Commander $\beta_{d}$ and $T_{d}$ maps (Planck Collaboration X, 2016) as the methodology employed both spatial and spectral information to disentangle the CIB contribution to the total emission (see discussion in Section 5.4). The GNILC spectral parameter fits show a large variability over the sky, when compared with the Commander spectral parameters, especially at high Galactic latitudes (Planck Collaboration Int. XLVIII, 2016). The informative priors on spectral parameters used in the Commander analysis restricts spatial variation of the best fit of spectral parameters in regions of low signal-to-noise, with a distribution that peaks close to the prior (Planck Collaboration X, 2016). The spatial variability seen in the GNILC spectral parameters correlates to structures seen in the dust intensity.
@@ 3.2.2 Dust Spectral Parameters
3.2.3 Synchrotron Spectral Parameters
@@
As described in Section 2.2.2, the synchrotron emission in all models developed here is governed by the spectral parameters $\beta_{s}$ and $c_{s}$. To build the large-scale template for the spatial variation of the synchrotron spectral index $\beta_{s}$, we begin with the full-sky $\beta_{s}$ map of Miville-Deschênes et al. (2008), obtained by combining the Haslam map in total intensity at 408 MHz (Remazeilles et al., 2015) and WMAP 3 yr K-band data (Hinshaw et al., 2007). The map has an angular resolution of about $7^{\circ}$ and was employed by previous PySM synchrotron models (Thorne et al., 2017).
@@
Incorporating newer constraints on synchrotron emission at 2.3 GHz from S-PASS, Krachmalnicoff et al. (2018) determined that this $\beta_{s}$ map underestimates the true level of spatial variations in $\beta_{s}$ across the Southern Galactic Hemisphere. Thus, we follow Krachmalnicoff et al. (2018) and rescale the $\beta_{s}$ map by first subtracting its mean value ($\bar{\beta}=-3.00$), multiplying the resulting map by a factor of $1.572$, and then adding back the mean value. This rescaling is further motivated by recent results in the Northern Galactic Hemisphere from QUIJOTE, which find even higher $\beta_{s}$ variation (de la Hoz et al., 2023; Rubiño-Martín et al., 2023).
@@
We next fit a power law to the power spectrum of this new template map over the multipole range $3610<\ell<$ 36, finding $\alpha_{\beta_{s}}=-0.61$, then construct a map from this power spectrum extrapolated to high multipoles. The $\beta_{s}$ small scales are modulated employing the same modulation map as the synchrotron intensity map (Section 3.1.2). As with the dust spectral parameter maps (Section 3.2.2), we combine this $-\ell$ map with the $-\ell$ template following the filter function of Equation 10.
@@
For the synchrotron curvature parameter $c_{s}$, there are no readily available template maps. The existing s3 model implements curvature as a sky-constant $c_{s}=-0.052$, consistent with the measurements from ARCADE ($c_{s}=-0.052\pm 0.005$, Kogut, 2012). To model a reasonable range of spatial variability, we assume that fluctuations in the value of $c_{s}$ follow the synchrotron intensity at large angular scales. Specifically, we start from the Haslam map 408 MHz $I$ map of Remazeilles et al. (2015), smoothed to a resolution of $5^{\circ}$. We then rescale this map with a linear transformation such that the (dimensionless) minimum and maximum pixel values over the full sky are $-0.076$ and $-0.041$ respectively, corresponding to the approximate range of $c_{s}$ measurements from ARCADE (Kogut, 2012, Figure 6). The resulting $c_{s}$ map has a mean and standard deviation of $-0.0517$ and $0.0054$, respectively, within the ARCADE footprint, in good agreement with the ARCADE measurements.
@@
To extend our $c_{s}$ map to angular scales $\ell>\ell_{1}=36$, we first generate a map of Gaussian random fluctuations with the same power law index as the $\beta_{s}$ map, i.e., $\alpha_{c_{s}}=\alpha_{\beta_{s}}=-0.61$. We modulate the resulting map with the Haslam 408 GHz $I$ map, smoothed to a $5^{\circ}$ FWHM resolution. We renormalize the map through a linear transformation such that the (dimensionless) pixel values range from 0.1–2. Finally, the modulated $-\ell$ map is combined with the $-\ell$ template following the filter function of Equation 10.
3.3 Summary of New Dust and Synchrotron Models
@@
d9 Modified blackbody GNILC PR2 $I$ + GNILC PR3 $Q$/$U$ 353 GHz Modulated + polarization fraction tensor Uniform $\beta_{d}$, $T_{d}$
d10 " " " $\beta_{d},\,T_{d}$ from GNILC PR2 Modulated
d11 " " " " " $I$, $Q$, $U$ $\beta_{d}$, $T_{d}$
d12 6 layers, each is a different modified blackbody GNILC PR2 $I$ + GNILC $Q$/$U$ 353 GHz Modulated + gaussian Random realization for each layer Random realization for each layer
@@
s4 Power law Haslam $I$ 408 MHz + WMAP $Q$/$U$ 23 GHz Modulated + polarization fraction tensor Uniform $\beta_{s}$
s5 " " " $\beta_{s}$ from Haslam, S-PASS, WMAP Modulated
s6 " " " " " $I$, $Q$, $U$ $\beta_{s}$
s7 Curved power law " " "+ $c_{s}$ from ARCADE "+ $c_{s}$ fluctuations
@@
The new dust and synchrotron models implemented here all improve on previous models through new data-driven templates, and use of the polarization fraction tensor framework to model small-scale fluctuations. Multiple models of dust and synchrotron are provided to explore a range of astrophysical complexity allowed by current constraints. In brief, the d9 and s4 models use the new data-driven templates and include $I$, $Q$, and $U$ fluctuations up to the largest supported $\ell$ values, but these models have sky-constant spectral parameters and thus no frequency decorrelation (see Section 5.3 for discussion). In contrast, d10 and s5 employ data-driven maps of spectral parameters to which small-scale fluctuations are added, inducing frequency decorrelation. The s7 synchrotron model provides an extra spectral parameter—curvature of the frequency spectrum—and thus additional complexity. The d11 and s6 models allow many realizations of d10 and s5, respectively, to be generated in which the large scales are fixed to the data-driven templates and the small scales have the same statistical properties but differing spatial morphologies. Tables 3 and 4 provide a high-level summary of these models. Extensive model comparisons are made in Section 5.
4 Other Models Implemented
4.1 Dust Layer Model
Evidence for variation of Galactic foreground emission laws as a function of frequency across the sky (e.g., Krachmalnicoff et al., 2018; Ade et al., 2025) implies that emission laws must also vary along the line of sight. As a consequence, even if dust emission can be described locally by a modified blackbody (Equation 1), a superposition of emission regions (i.e., an integral along the line of sight) is not a modified blackbody. In addition, if along the line of sight different line elements emit polarized radiation with different polarization angles, the frequency scaling can vary between $I$, $Q$ and $U$ (Tassis & Pavlidou, 2015). Evidence for this “line-of-sight frequency decorrelation” has been found in Planck data (Pelgrims et al., 2021).
To model the complexity of multiple layers of dust along the line of sight, we base our PySM 3 implementation d12 on the approach of Martínez-Solaeche et al. (2018) in the PSM software [^11] See version 2.3.3 https://apc.u-paris.fr/~delabrou/PSM/psm.html..
The PSM was used to produce six maps of dust emission $S_{\nu_{0}}^{k}(\theta)$ at $\nu_{0}=353$ GHz (intensity and polarization, for a total of 18 HEALPix maps at $N_{\rm side}$=2048), and six maps of dust spectral index $\beta_{d}^{k}(\theta)$ and dust temperature $T_{d}^{k}(\theta)$ at $N_{\rm side}$=2048, following the approach described in Martínez-Solaeche et al. (2018), to form six different layers.
@@ 4.1 Dust Layer Model
where $S_{\nu}(\theta)$ stands for any of the three Stokes parameters of interest, $I$, $Q$ and $U$, at frequency $\nu$ at sky location $\theta$, and superscripts ${(k)}$ indicate the layer, from one to six.
In the implementation used here, the 353 GHz templates for the six emission layers are slightly different from those of Martínez-Solaeche et al. (2018). They are generated at HEALPix $N_{\rm side}=2048$ instead of 512, with a revised version of the pipeline. Intensity maps are based on the Planck PR2 GNILC maps (Planck Collaboration Int. XLVIII, 2016), while large scale polarized emission maps are the GNILC-derived maps from Martínez-Solaeche et al. (2018). These are complemented by small-scale realizations with scale dependence matching the $TT$, $EE$ and $BB$ dust spectra measured in Planck Collaboration XI (2020), modulated by the large-scale local intensity and polarized intensity level. Specifically, for each layer $k$, and for each of $T$, $E$ and $B$, we model the final emission in harmonic space as:
@@
where $X^{(k)}_{\nu_{0}}$ is the observed emission deconvolved from the instrumental beam, $Y^{(k)}_{\nu_{0}}$ is a randomly generated set of harmonic coefficients with harmonic spectrum following Planck constraints (Planck Collaboration XI, 2020), and $h_{\ell}$ is a window defining the transition between the two regimes. For intensity, $h_{\ell}$ corresponds to a $5\arcmin$ beam window function, while for polarization, we use a $150\arcmin$ beam for the three nearest layers (which dominate at high Galactic latitude) and a $120\arcmin$ beam for the three farthest layers (which dominate near and in the Galactic plane).
@@
In this model, maps of spectral parameters ($\beta_{d}$ and $T_{d}$) are randomly generated for each layer as described in Martínez-Solaeche et al. (2018). As a consequence, they are not constrained to match the observed temperature and spectral index maps. While their statistics (e.g., amplitude, correlation between $\beta_{d}$ and $T_{d}$) are compatible with those of real data, they provide a different realization, which results in increased differences between the real sky and the model at frequencies increasingly far from the reference frequency, $\nu_{0}=353$ GHz.
4.2 CO Models

### v2

@@ III.1.4 Synchrotron Amplitude
Deriving a full-sky template for synchrotron emission in total intensity is challenging, both due to the paucity of low-frequency surveys with full sky coverage and the difficulty in disentangling synchrotron emission from free-free emission and AME at frequencies above $\sim$10 GHz. We follow Thorne:2017 in basing our template on the Haslam 408 MHz survey (Haslam:1982) and in particular the reprocessing of the original maps by Remazeilles:2015. We rescale the 408 MHz map to 23 GHz assuming a sky-constant scaling of the brightness temperature with $\nu^{-3.1}$. Finally, we smooth the resulting map with a $2^{\circ}$ Gaussian beam.
@@
A key limitation of the current synchrotron total intensity template is that, despite the point source masking, we still find excess power from Galactic point sources at low Galactic latitudes near the Galactic center. This excess power can then leak into our final polarization maps through the polarization fraction tensor transformations (Equations (III.1.1) and (III.1.1)). Our models are unaffected by this issue at $\ell\lesssim 3000$ and outside the Galactic plane. Addressing small-scale structures in bright regions will be deferred to a future release.
@@
III.2 Small-Scale Fluctuations in Spectral Parameters
III.2.1 Methods Overview
@@
As there is no sense of polarization in the spectral parameter maps, we do not utilize the polarization fraction tensor framework, but rather work with standard power spectra. Given a template map $\mathcal{T}$ of some spectral parameter $X$, we first fit the power spectrum of $\mathcal{T}$ with a power law $\mathcal{D}_{\ell}^{XX}\propto\ell^{\alpha}$ over a range of scales $\ell_{0}<\ell<\ell_{1}$ where it is well-measured. We then generate a map of small-scale fluctuations using the power law fit. We multiply the resulting map by a modulation map, in analogy with the amplitude modulations described in Section III.1.2. Where there is more dust there is likely to be both more discrete environments (either within a beam or along the line of sight) and more variation of dust properties within a region (e.g., the interior versus exterior of a molecular cloud, a photo-dissociation region, etc.).
Finally, we combine the template map with the synthesized small scales using the filter function in Equation 10. As with the amplitude parameters, we use the input template maps for the spectral parameters as-is in the inner Galactic plane, defined by the Planck GAL097 mask.
@@
III.2.2 Dust Spectral Parameters
As described in Section II.2.1, the dust emission in all models developed here is governed by the spectral parameters $\beta_{d}$ and $T_{d}$. For the $\beta_{d}$ and $T_{d}$ template maps, we employ the $\beta_{d}$ and $T_{d}$ maps [^10] COM_CompMap_Dust-GNILC-Model-Spectral-Index_2048_R2.00.fits, COM_CompMap_Dust-GNILC-Model-Temperature_2048_R2.00.fits derived from 2015 Planck GNILC component separation analysis (planck2016-XLVIII). These maps have lower CIB contamination than the Commander $\beta_{d}$ and $T_{d}$ maps (planck2014-a12) as the methodology employed both spatial and spectral information to disentangle the CIB contribution to the total emission (see discussion in Section V.4). The GNILC spectral parameter fits show a large variability over the sky, when compared with the Commander spectral parameters, especially at high Galactic latitudes (planck2016-XLVIII). The informative priors on spectral parameters used in the Commander analysis restricts spatial variation of the best fit of spectral parameters in regions of low signal-to-noise, with a distribution that peaks close to the prior (planck2014-a12). The spatial variability seen in the GNILC spectral parameters correlates to structures seen in the dust intensity.
@@ III.2.2 Dust Spectral Parameters
III.2.3 Synchrotron Spectral Parameters
@@
As described in Section II.2.2, the synchrotron emission in all models developed here is governed by the spectral parameters $\beta_{s}$ and $c_{s}$. To build the large-scale template for the spatial variation of the synchrotron spectral index $\beta_{s}$, we begin with the full-sky $\beta_{s}$ map of Miville-Deschenes:2008, obtained by combining the Haslam map in total intensity at 408 MHz (Haslam:1982) and WMAP 3 yr K-band data (Hinshaw:2007). The map has an angular resolution of about $7^{\circ}$ and was employed by previous PySM synchrotron models (Thorne:2017).
@@
Incorporating newer constraints on synchrotron emission at 2.3 GHz from S-PASS, Krachmalnicoff:2018 determined that this $\beta_{s}$ map underestimates the true level of spatial variations in $\beta_{s}$ across the Southern Galactic Hemisphere. Thus, we follow Krachmalnicoff:2018 and rescale the $\beta_{s}$ map by first subtracting its mean value ($\bar{\beta}=-3.00$), multiplying the resulting map by a factor of $1.572$, and then adding back the mean value. This rescaling is further motivated by recent results in the Northern Galactic Hemisphere from QUIJOTE, which find even higher $\beta_{s}$ variation (delaHoz:2023; Rubino-Martin:2023).
@@
We next fit a power law to the power spectrum of this new template map over the multipole range $10<\ell<36$, finding $\alpha_{\beta_{s}}=-0.61$, then construct a map from this power spectrum extrapolated to high multipoles. The $\beta_{s}$ small scales are modulated employing the same modulation map as the synchrotron intensity map (Section III.1.2). As with the dust spectral parameter maps (Section III.2.2), we combine this high-$\ell$ map with the low-$\ell$ template following the filter function of Equation 10.
@@
For the synchrotron curvature parameter $c_{s}$, there are no readily available template maps. The existing s3 model implements curvature as a sky-constant $c_{s}=-0.052$, consistent with the measurements from ARCADE ($c_{s}=-0.052\pm 0.005$, Kogut:2012). To model a reasonable range of spatial variability, we assume that fluctuations in the value of $c_{s}$ follow the synchrotron intensity at large angular scales. Specifically, we start from the Haslam map 408 MHz $I$ map of Remazeilles:2015, smoothed to a resolution of $5^{\circ}$. We then rescale this map with a linear transformation such that the (dimensionless) minimum and maximum pixel values over the full sky are 0.076-0.076 and 0.041-0.041 respectively, corresponding to the approximate range of $c_{s}$ measurements from ARCADE (Kogut:2012, Figure 6). The resulting $c_{s}$ map has a mean and standard deviation of 0.0517-0.0517 and $0.0054$, respectively, within the ARCADE footprint, in good agreement with the ARCADE measurements.
@@
To extend our $c_{s}$ map to angular scales $\ell>\ell_{1}=36$, we first generate a map of Gaussian random fluctuations with the same power law index as the $\beta_{s}$ map, i.e., $\alpha_{c_{s}}=\alpha_{\beta_{s}}=-0.61$. We modulate the resulting map with the Haslam 408 GHz $I$ map, smoothed to a $5^{\circ}$ FWHM resolution. We renormalize the map through a linear transformation such that the (dimensionless) pixel values range from 0.1–2. Finally, the modulated high-$\ell$ map is combined with the low-$\ell$ template following the filter function of Equation 10.
III.3 Summary of New Dust and Synchrotron Models
@@
d9 Modified blackbody
GNILC PR2 II +
GNILC PR3 QQ/$U$
353 GHz
Modulated +
polarization
fraction tensor
Uniform $\beta_{d}$, $T_{d}$ Uniform $\beta_{d}$, $T_{d}$
d10 " " "
$\beta_{d},\,T_{d}$ from
GNILC PR2
Modulated
d11 " " " " "
II, $Q$, $U$ &
$\beta_{d}$, $T_{d}$
d12
Six layers, each with a
different modified
blackbody
GNILC PR2 II +
GNILC QQ/$U$
353 GHz
Modulated +
Gaussian
Random realization
for each layer
Random realization
for each layer
@@
s4 Power law
Haslam $I$ 408 MHz +
WMAP QQ/$U$ 23 GHz
Modulated +
polarization
fraction tensor
Uniform $\beta_{s}$ Uniform $\beta_{s}$
s5 " " "
$\beta_{s}$ from Haslam,
S-PASS, WMAP
Modulated
s6 " " " " " II, $Q$, $U$ & $\beta_{s}$
s7 Curved power law " "
"+ $c_{s}$
from ARCADE
"+ $c_{s}$
fluctuations
@@
The new dust and synchrotron models implemented here all improve on previous models through new data-driven templates, and use of the polarization fraction tensor framework to model small-scale fluctuations. Multiple models of dust and synchrotron are provided to explore a range of astrophysical complexity allowed by current constraints. In brief, the d9 and s4 models use the new data-driven templates and include $I$, $Q$, and $U$ fluctuations up to the largest supported $\ell$ values, but these models have sky-constant spectral parameters and thus no frequency decorrelation (see Section V.3 for discussion). In contrast, d10 and s5 employ data-driven maps of spectral parameters to which small-scale fluctuations are added, inducing frequency decorrelation. The s7 synchrotron model provides an extra spectral parameter—curvature of the frequency spectrum—and thus additional complexity. The d11 and s6 models allow many realizations of d10 and s5, respectively, to be generated in which the large scales are fixed to the data-driven templates and the small scales have the same statistical properties but differing spatial morphologies. Tables 3 and 4 provide a high-level summary of these models. Extensive model comparisons are made in Section V.
IV Other Models Implemented
In this section, we discuss the construction of other new foreground models, including the dust d12 and CO models (co1, co2, and co3), that employ different methods from those outlined in Section III.
IV.1 Dust Layer Model
Evidence for variation of Galactic foreground emission laws as a function of frequency across the sky (e.g., Krachmalnicoff:2018; Ade:2025b) implies that emission laws must also vary along the line of sight. As a consequence, even if dust emission can be described locally by a modified blackbody (Equation 1), a superposition of emission regions (i.e., an integral along the line of sight) is not a modified blackbody. In addition, if along the line of sight different line elements emit polarized radiation with different polarization angles, the frequency scaling can vary between $I$, $Q$ and $U$ (Tassis:2015). Evidence for this “line-of-sight frequency decorrelation” has been found in Planck data (Pelgrims:2021).
To model the complexity of multiple layers of dust along the line of sight, we base our PySM 3 implementation d12 on the approach of Martinez-Solaeche:2018 in the PSM software [^11] See version 2.3.3 https://apc.u-paris.fr/~delabrou/PSM/psm.html..
The PSM was used to produce six maps of dust emission $S_{\nu_{0}}^{k}(\theta)$ at $\nu_{0}=353$ GHz (intensity and polarization, for a total of 18 HEALPix maps at $N_{\rm side}$=2048). It was also used to generate six maps of dust spectral index $\beta_{d}^{k}(\theta)$ and dust temperature $T_{d}^{k}(\theta)$ at $N_{\rm side}$=2048, one per layer, following the approach described in Martinez-Solaeche:2018. This set of maps is used to form six different layers of dust emission.
@@ IV.1 Dust Layer Model
where $S_{\nu}(\theta)$ stands for any of the three Stokes parameters of interest, $I$, $Q$ and $U$, at frequency $\nu$ at sky location $\theta$, and superscripts (k){(k)} indicate the layer, from one to six.
In the implementation used here, the 353 GHz templates for the six emission layers are slightly different from those of Martinez-Solaeche:2018. They are generated at HEALPix $N_{\rm side}=2048$ instead of 512, with an updated version of the Martinez-Solaeche:2018 pipeline that supports different $N_{\rm side}$ values. Intensity maps are based on the Planck PR2 GNILC maps (planck2016-XLVIII), while large scale polarized emission maps are the GNILC-derived maps from Martinez-Solaeche:2018. These are complemented by small-scale realizations with scale dependence matching the $TT$, EEEE and $BB$ dust spectra measured in planck2016-l11A, modulated by the large-scale local intensity and polarized intensity level. Specifically, for each layer $k$, and for each of $T$, $E$ and $B$, we model the final emission in harmonic space as:
@@
where $X^{(k)}_{\nu_{0}}$ is the observed emission deconvolved from the instrumental beam, $Y^{(k)}_{\nu_{0}}$ is a randomly generated set of harmonic coefficients with harmonic spectrum following Planck constraints (planck2016-l11A), and $h_{\ell}$ is a window defining the transition between the two regimes. For intensity, $h_{\ell}$ corresponds to a $5\arcmin$ beam window function, while for polarization, we use a $150\arcmin$ beam for the three nearest layers (which dominate at high Galactic latitude) and a $120\arcmin$ beam for the three farthest layers (which dominate near and in the Galactic plane).
@@
In this model, maps of spectral parameters ($\beta_{d}$ and $T_{d}$) are randomly generated for each layer as described in Martinez-Solaeche:2018. As a consequence, they are not constrained to match the observed temperature and spectral index maps. While their statistics (e.g., amplitude, correlation between $\beta_{d}$ and $T_{d}$) are compatible with those of real data, they provide a different realization, which results in increased differences between the real sky and the model at frequencies increasingly far from the reference frequency, $\nu_{0}=353$ GHz.
IV.2 CO Models